### v1

# Efficient Algorithm Level Error Detection for Number-Theoretic Transform Assessed on FPGAs

Abstract
Polynomial multiplication stands out as a highly demanding arithmetic process in the development of post-quantum cryptosystems. The importance of number-theoretic transform (NTT) extends beyond post-quantum cryptosystems, proving valuable in enhancing existing security protocols such as digital signature schemes and hash functions. Due to the potential for errors to significantly disrupt the operation of secure, cryptographically-protected systems, compromising data integrity, and safeguarding against side-channel attacks initiated through faults it is essential to incorporate mitigating error detection schemes. This paper introduces algorithm level fault detection schemes in NTT multiplication, representing a significant enhancement compared to previous research. We evaluate this through the simulation of a fault model, ensuring that the conducted assessments accurately mirror the obtained results. Consequently, we attain a notably comprehensive coverage of errors. Finally, we assess the performance of our efficient error detection scheme on FPGAs to showcase its implementation and resource requirements. Through implementation of our error detection approach on Xilinx/AMD Zynq Ultrascale+ and Artix-7, we achieve a comparable throughput with just a 9% increase in area and 13% increase in latency compared to the original hardware implementations.
Index Terms:
Fault detection, fast Fourier transform (FFT), FPGA, number-theoretic transform (NTT).
I Introduction
Fast Fourier Transform (FFT) [1] algorithms which are used to compute the Discrete Fourier Transform (DFT) have various applications, spanning from digital signal processing to the efficient multiplication of large integers. When the coefficients of the polynomial are specifically chosen from a finite field, the resulting transform is known as the Number Theoretic Transform (NTT) [2], and it can be computed using FFT algorithms designed for operations within this particular finite field. An efficient approach for polynomial multiplication, NTT holds significant importance in post-quantum cryptosystems, e.g., lattice-based cryptosystems which is regarded as a leading contender for quantum-secure public key cryptography, primarily because of its broad applicability and security proofs grounded in the worst-case hardness of established lattice problems. The NTT has proven to be a potent tool that facilitates the computation of this operation with quasi-polynomial complexity $O(n.lgn)$. Several recent studies have been conducted on optimizing NTT [3], [4], [5], [6], [7], [8], and [9]. In addition to polynomial multiplication, the use of NTT has the potential to significantly enhance existing schemes by improving their security parameters. NTT is widely used in signature schemes, hash functions, and identification schemes. As a result, incorporating efficient error detection mechanisms in NTT for polynomial multiplication will not only improve the security and reliability but also mitigate the risk of fault attacks in the respective algorithms in post-quantum cryptography. Moreover, such schemes could at least alleviate the attack surface or be add-ons to other approaches.
I-A Related works
We catogorize relevant studies into three subsets. The initial category centers on fault attacks and countermeasures applied to post-quantum cryptographic schemes, particularly those extensively utilizing NTT. Ravi et al. [10] presented the first fault injection analysis of the NTT. The authors found a significant flaw in the way NTT is implemented in the pqm4 library [11]. The identified vulnerability, referred to as twiddle-pointer, is exploited to demonstrate practical and effective attacks on Kyber and Dilithium [12]. Several other authors have reported employing fault injection on structured lattice-based schemes as a foundation for attacks [13], [14], [15], and [16] threatening the security of implementations.
The focus of the second category lies in offering fault detection methodologies pertaining to the algorithmic level of classical cryptographic schemes. In [17], the authors suggested an effective algorithmic-level error detection for the ECSM window method, aiming to identify both permanent and transient errors. In [18], the authors presented error detection schemes designed for the double-and-add-always and Montgomery ladder ECSM algorithm utilized in non-supersingular elliptic curves.
The third category focuses on developing fault detection schemes specifically for NTT. In [19], the authors integrate the advantages of bitslicing, a software implementation technique where a datapath of an $n$-bit processor is treated as $n$ parallel single-bit datapaths, to devise a fault countermeasure for the NTT used in Dilithium [20]. Sarker et al. [21] and [22], presented error detection architectures of NTT based on recomputation with encoded operands. The authors achieved high error coverage with low area overhead. Despite the recomputation mechanism, the latency is elevated. We note that despite high error coverage, for fast implementations, these works might increase the total time to levels not acceptable.
I-B Our Major Contributions
  • We have proposed an algorithm level fault detection scheme of NTT multiplication which is widely used in lattice based cryptography. We achieved high error coverage with less latency compared to previous works. This was achieved through the implementation of algorithm-level error detection for NTT section and partial recomputation for pre-computation sub-block. We performed simulations for single and burst fault injection using our proposed scheme. The simulation outcomes demonstrated that our approach is capable of detecting diverse types of faults with high error coverage and offers protection against the NTT fault attack presented in [10] and those with respective fault models.
  • Our error detection method is deployed on Xilinx/AMD Zynq Ultrascale+ and Artix-7. The results of our implementations indicate that we can attain a significantly high level of error coverage with only a 13% increase in latency and a 9% area bloat.
II Preliminaries
II-A Number Theoretic Transform (NTT)
[…]
The NTT of a sequence $a$ is derived as:
$\widetilde{a}=NTT(a)=\left[\begin{array}[]{cccc}\omega^{0}&\omega^{0}&...&\omega^{0}\\ \omega^{0}&\omega^{1}&...&\omega^{n-1}\\ \omega^{0}&\omega^{2}&...&\omega^{2(n-1)}\\ .&.&...&.\\ \omega^{0}&\omega^{n-1}&...&\omega^{(n-1)^{2}}\end{array}\right]*\left[\begin{array}[]{c}a(0)\\ a(1)\\ a(2)\\ .\\ a(n-1)\end{array}\right]$ (1)
[…]
Algorithm 1 Iterative NTT [24]
Input: $a\in\mathbb{Z}_{q}$ of length $n=2^{k}$, $\omega$ is primitive $n$-th root of unity
Output: $\widetilde{a}=\text{NTT}(a)$ in bit-reversed order
1: $\widetilde{a}=a$
2: for $s=k$ to $1$:
3:       $m=2^{s}$
4:       for $k=0$ to $12^{k-s}-$ 1:
5:              $\omega=\omega^{\text{bit-reverse}(k)\,.\,m/2}$
6:              for $j=0$ to $m/2-1$:
7:                    $u=\widetilde{a}[k.m+j]$
8:                    $t\equiv\omega\,.\,\widetilde{a}[k.m+j+m/2]$ mod $q$
9:                    $\widetilde{a}[k.m+j]\equiv(u+t)$ mod $q$
10:                  $\widetilde{a}[k.m+j+m/2]\equiv(u-t)$ mod $q$
11: return $\widetilde{a}$
II-B Negative Wrapped Convolution
[…]
Algorithm 2 Pre-processing step for NTT calculations
[…]
Output: $\hat{a}=\text{pre-process}(a)$
1: $\hat{a}=a$
2: for $i=0$ to $n-1$:
3:       $\hat{a}[i]=a[i]\,.\,$$\psi^{i}$
4: return $\hat{a}$
III Proposed Error Detection Scheme
In this section, we present our error detection schemes on polynomial multiplication, $\text{NTT}(\hat{a})\,\Circle\,\text{NTT}(\hat{b})$ where $\hat{a}=\text{pre-process}(a)$ and $\hat{b}=\text{pre-process}(b)$. We have devised an error detection scheme at the algorithm level on the component-wise NTT multiplication sub-block, $\text{NTT}(\hat{a})\,\Circle\,\text{NTT}(\hat{b})$, as well as an error detection scheme for the pre-process function, $\hat{a}=\text{pre-process}(a)$, involving careful utilization of recomputation using shifted operands.
Figure 1: Concurrent error detection scheme for NTT operation which is used in [25].
III-1 Component-wise NTT Multiplication
Our focus is on $\text{NTT}(\hat{a})\,\Circle\,\text{NTT}(\hat{b})$ operation in this section. Jou et al. [25], proposed an algorithm level error detection scheme on FFT network. Their proposed error detection scheme is depicted in Fig. 1. In order to achieve an efficient scheme, we have adapted the aforementioned approach to component-wise NTT multiplication. The primary focus is to use an error detection scheme with small overhead and high error detection ratio with respect to efficient realizations of the NTT. From NTT definition, we can use the relationships presented below:
$\widetilde{a}(0)=\sum_{j=0}^{n-1}a(j)\omega^{0}=\sum_{j=0}^{n-1}a(j),\quad j=0,1,...,n-1$ (2)
$\widetilde{b}(0)=\sum_{j=0}^{n-1}b(j)\omega^{0}=\sum_{j=0}^{n-1}b(j),\quad j=0,1,...,n-1$ (3)
The result of the NTT multiplication for index 0 can be written as:
$\widetilde{a}(0)\widetilde{b}(0)=\sum_{j=0}^{n-1}a(j)\sum_{j=0}^{n-1}b(j)$ (4)
We note that (4) is used in our proposed error detection scheme.
If we rotate the input sequence of (1) by one, then we get:
$\left[\begin{array}[]{c}\widetilde{a}(0)\\ \widetilde{a}(1)\\ \widetilde{a}(2)\\ .\\ \widetilde{a}(N-1)\end{array}\right]$ $=\rho*\left[\begin{array}[]{c}\hat{a}(1)\\ \hat{a}(2)\\ .\\ \hat{a}(N-1)\\ \hat{a}(0)\end{array}\right]$ (15)
or
$\left[\begin{array}[]{c}\widetilde{a}(0)/\omega^{0}\\ \widetilde{a}(1)/\omega^{1}\\ \widetilde{a}(2)/\omega^{2}\\ .\\ \widetilde{a}(n-1)/\omega^{(n-1)}\end{array}\right]$ $=\theta*\left[\begin{array}[]{c}\hat{a}(1)\\ \hat{a}(2)\\ .\\ \hat{a}(n-1)\\ \hat{a}(0)\end{array}\right]$ (26)
where the symbol * denotes matrix multiplication, $\rho=\left[\begin{array}[]{cccc}\omega^{0}&\omega^{0}&...&\omega^{0}\\ \omega^{1}&\omega^{2}&...&\omega^{0}\\ \omega^{2}&\omega^{4}&...&\omega^{0}\\ .&.&...&.\\ \omega^{n-1}&\omega^{2(n-1)}&...&\omega^{0}\end{array}\right]$ in (5), and $\theta=$$\left[\begin{array}[]{cccc}\omega^{0}&\omega^{0}&...&\omega^{0}\\ \omega^{0}&\omega^{1}&...&\omega^{n-1}\\ \omega^{0}&\omega^{2}&...&\omega^{2(n-1)}\\ .&.&...&.\\ \omega^{0}&\omega^{n-1}&...&\omega^{(n-1)^{2}}\end{array}\right]$ in (6). From $\alpha$ (in (1)) + $\beta$ (in (6)), we get the following:
$\left[\begin{array}[]{c}\widetilde{a}(0)(\alpha+\beta\omega^{-0})\\ \widetilde{a}(1)(\alpha+\beta\omega^{-1})\\ \widetilde{a}(2)(\alpha+\beta\omega^{-2})\\ .\\ \widetilde{a}(n-1)(\alpha+\beta\omega^{-(n-1)})\end{array}\right]$ $=\theta*\left[\begin{array}[]{c}\alpha\hat{a}(0)+\beta\hat{a}(1)\\ \alpha\hat{a}(1)+\beta\hat{a}(2)\\ .\\ \alpha\hat{a}(n-2)+\beta\hat{a}(n-1)\\ \alpha\hat{a}(n-1)+\beta\hat{a}(0)\end{array}\right]$ (37)
where $\alpha$ and $\beta$ in (7) are scalars which can be selected arbitrarily. Without losing generality and for the sake of simplicity, we select $\alpha=2$ and $\beta=1$. Let us denote $\hat{a}^{m}$ and $\hat{b}^{m}$, respectively, as the inputs which are shifted by $m$. In other words, we can achieve:
$\widetilde{a}=\frac{1}{(\alpha+\beta\omega^{-k})}NTT(\alpha\hat{a}+\beta\hat{a}^{1})$ (38)
$\widetilde{b}=\frac{1}{(\alpha+\beta\omega^{-k})}NTT(\alpha\hat{b}+\beta\hat{b}^{1})$ (39)
Let us denote $c$ as the component-wise NTT multiplication of $\hat{a}$ and $\hat{b}$. We can get:
$c$ $=NTT(\hat{a})\,\Circle\,NTT(\hat{b})=$
$\frac{1}{(\alpha+\beta\omega^{-k})^{2}}NTT(\alpha\hat{a}+\beta\hat{a}^{1})\,\Circle\,NTT(\alpha\hat{b}+\beta\hat{b}^{1})$ (40)
The proposed error detection scheme is depicted in Fig. 2. The Decoder_2 sub-block multiplies its input with $\frac{1}{(\alpha+\beta\omega^{-k})^{2}}$, which is the inverse of $(\alpha+\beta\omega^{-k})^{2}$ in $\mathbb{R}=\mathbb{Z}[X]/(X^{n}+1)$. The scheme for error detection involves comparing $c(0)$, obtained from (10), with $\sum_{j=0}^{n-1}a(j)\sum_{j=0}^{n-1}b(j)$.
Figure 2: Proposed algorithm level error detection scheme for NTT multiplication module.
III-2 Pre-processor
The pre-processor module performs element-by-element multiplication on two input arrays. In proposing the error detection scheme, we have applied an additional recomputing with shifted operands. As shown in Fig. 3, the encoding and decoding modules constitute shifting which is free in hardware. The pre-processing module considers a slight overhead linked to NTT multiplication. As a result, the error detection scheme employing recomputation does not notably impact the overall performance.
Figure 3: Proposed error detection scheme for pre-process module through recompuation with shifted operands.
IV Simulations and Fault Model
In this section, we present the utilized fault model and the simulation results acquired by implementing Python code to show error detection coverage, employing the presented fault model, and utilizing our efficient error detection schemes in this paper.
IV-A Fault Model
We assume that faults occur in either the complex multiplier or adder, or a specific group of input or output lines in the butterfly module which is used in NTT. As depicted in Fig. 4, faults can happen at modules $\{1,2,3\}$ marked in the figure, which may result in erroneous outputs of a butterfly sub-block used in NTT module. Fault occurrence follow a normal distribution within butterfly modules in Fig. 2. Furthermore, as we provide fault detection schemes over NTT multiplication, faults can happen during component wise multiplication module.
Figure 4: The utilized fault model in this work for the butterfly sub-block.
IV-B Error Coverage Simulation Results
To assess the error coverage, we employed Python3 to perform simulations, applying the described fault models and error detection techniques to both the NTT and pre-process modules. The simulations encompassed a million samples, using parameters identical to the standard Kyber $(n=256)$ and BLISS signature [26] $(n=512)$. As shown in Table I, with increase in the number of faults, we can attain an error detection ratio close to 100% for both modules. Higher error coverage is achievable for identifying burst errors. We have done extensive experiments injecting burst adjacent faults, the results of which are not presented for the sake of brevity.
Table I: Achieved error detection ratios of the proposed schemes with 1, 2, 4, 8, and 16 faults occurrence for one million samples
Parameters ${}^{1}n=256,$${}^{2}\omega=17$, $n=512,\omega=2,401$,
${}^{3}q=3,329$ $q=12,289$
Component Pre-process NTT Pre-process NTT
Mult Mult
1 99.6% 51% 99.7% 53%
Number of 2 99.9% 67% 99.9% 70.6%
faults 4 100% 90.3% 100% 90.9%
8 100% 99% 100% 99%
16 100% 99.9% 100% 99.9%
${}^{1}$$n=$ Polynomial size, ${}^{2}$$\omega=$ Twiddle factor, ${}^{3}$$q=$ Prime number
V Xilinx/AMD FPGA Implementation results
To confirm the efficiency of our presented approach, we chose to assess its performance by applying it to the NTT multiplication involving a 256-degree polynomial. We conducted a benchmark for implementation on different FPGAs, i.e., Xilinx/AMD Zynq Ultrascale+ and Artix-7. The results clearly demonstrate that our proposed efficient error detection schemes maintain a modest overhead while effectively achieving a high level of fault detection. We utilized the High-Level Synthesis (HLS) Vitis development environment to transform our proposed schemes into Register Transfer Level (RTL) hardware descriptions. The generated IP imported to Vivado to report power, utilization, and latency. Tables II and III demonstrate the results of our implementations and the derivations for area, timing, power, and energy. In the Vivado synthesis tool context, area is determined by combining Slices and DSPs, with a conversion ratio where one DSP is deemed equivalent to 100 Slices. Our proposed error detection designs achieve a maximum operational frequency of around 140 MHz across all the FPGAs. Our proposed error detection scheme imposes a maximum overhead of 9% in additional area and introduces a latency increase of up to 13% in clock cycles, at most considering different designs. Our simulation and implementation code is publicly available in our github [^1] https://github.com/KasraAhmadi/NTT_Error_Detection.
V-A Implementation Optimization
Although HLS has shifted the design entry level of abstraction from RTL to C/C++, practical implementation often requires significant source code rewriting to make it HLD-ready. We have carefully taken this into effort as discussed here. This includes the incorporation of pragmas to attain satisfactory performance. Our intended area and timing efforts implementation was realized by strategically inserting the following pragmas into our program.
Table II: AMD/Xilinx Zynq Ultrascale+, xczu4ev-sfvc784-2-i Implementation Results
Strategy Area Timing
Effort Effort
Scheme Our Baseline Our Baseline
scheme work scheme work
LUTs 1,469 1,435 3,207 3,165
Area FFs 1,188 1,122 2,227 2,161
CLBs 343 340 733 683
DSPs 33 30 54 51
Power (W) 0.46 0.45 0.52 0.51
@ 140 MHz
Latency 3,703 3,438 2,241 1,979
Timing [CCs]
Total time 26,439 24,547 16,000 14,130
[ns]
Energy (nJ) 12,161 11,046 8,320 7,206
Our Scheme: The design which includes error detection scheme.
Baseline work: The design which does not include any error detection scheme.
Table III: AMD/Xilinx Artix-7, xc7a75ti-ftg256-1L Implementation Results
Strategy Area Timing
Effort Effort
Scheme Our Baseline Our Baseline
scheme work scheme work
LUTs 1,569 1,530 2,956 2,939
Area FFs 1,693 1,597 2,979 2,883
SLICEs 673 656 1,240 1,253
DSPs 33 30 54 51
Power (W) 0.2 0.19 0.28 0.26
@ 140 MHz
Latency 3,749 3,482 2,267 2,003
Timing [CCs]
Total time 26,767 24,861 16,186 14,301
[ns]
Energy (nJ) 5,353 4,723 4,532 3,718
V-A1 Pre-calculate the decoder module values
To enhance the efficiency of the proposed error detection scheme, we have the option to pre-calculate the decoder module in (10), specifically computing $\frac{1}{(\alpha+\beta\omega^{-k})^{2}}$ for various values of the parameter $k$ and save them in memory.
V-A2 Utilizing task-level pipelining
We perform loop unrolling on the outer loop in Algorithm 1 (Line 2) and transform the remaining code into a function named "stage," illustrating each stage in the NTT architecture. We undertook this approach to leverage task-level pipelining, allowing functions and loops to operate simultaneously. This enhances the concurrency of the RTL implementation, resulting in an overall increase in design throughput.
V-A3 Pipelining the stage function
This pragma was inserted to facilitate instruction-level pipelining, aiming to enhance throughput and clock frequency within each NTT stage function. It is important to acknowledge that this optimization entails the trade-off of utilizing extra digital resources.
VI Conclusion
In this paper, we proposed an efficient algorithm level error detection scheme over polynomial multiplication using NTT. We have aimed for low hardware overhead and low latency, suitable for deeply-embedded systems. The presented scheme effectively safeguards cryptographic algorithms that employ NTT multiplication. By performing simulations, we demonstrated that the proposed error detection scheme achieved extensive coverage of faults. Moreover, we implemented our proposed design on two Xilinx/AMD FPGAs. Regarding overhead and latency, the implementation led to minimal additional expenses in hardware (at most 9% area, 13% latency, 7% power, and 13% energy). With high error coverage and acceptable overhead, the proposed schemes in this works are suitable for resource-constrained and sensitive usage models.
Acknowledgments
This work was supported by the US National Science Foundation (NSF) through the award SaTC-1801488.